# Folded Context Condensation in Path Integral Formalism for Infinite Context Transformers

### v4

@@ I Introduction
Transformers architecture [1] has revolutionized natural language processing (NLP) by demonstrating unparalleled effectiveness in capturing long-range dependencies and enabling scalable, parallelized computation. Unlike earlier Recurrent Neural Network (RNN) based models, which process sequences sequentially and suffer from vanishing gradient issues [2], Transformers leverage self-attention mechanisms to establish global dependencies across tokens, leading to significant improvements in contextual reasoning and coherent text generation. This architectural advantage has propelled Transformers to the forefront of modern NLP, underpinning the development of large-scale language models (LLMs) with remarkable generative capabilities.
Despite their empirical success, the theoretical foundations of Transformer architectures remain incomplete. While extensive research has explored their scaling properties [3], emergent behaviors [4], and empirical performance [5, 6], these studies provide limited insight into the fundamental mechanisms governing Transformer architectures. In particular, there is a lack of rigorous theoretical formulations explaining how information propagates through layers and how hierarchical structures are captured within Transformer models. This gap in understanding limits the interpretability and optimization of Transformer-based architectures and presents challenges in extending their capabilities systematically.
Another critical limitation of Transformers is their computational inefficiency when handling long sequences. The self-attention mechanism inherently scales quadratically with sequence length in both memory and computation, making it impractical to process extensive contexts within a single pass. This limitation hinders the retention of dependencies across long documents, dialogues, or continuous interactions, prompting the development of alternative strategies such as Infini-Attention [7], which restructures the self-attention mechanism by introducing a hybrid memory system that allows interaction between local context tokens and external memory components, leveraging a linear attention [8] framework to reduce computational overhead. Another innovative solution is Titans [9], which incorporates trainable neural long-term memory, enabling the model to retain and retrieve past contextual information beyond traditional Transformer constraints. However, these methods often introduce external memory components that do not engage with the Feed-Forward Network, which we believe it may reduce their ability to convey context information across segments.
@@
Building on this theoretical foundation in Section II, we present an implementation in Section III and IV that extends Transformer architectures with a time-evolving memory buffer. This structured approach ensures that past token representations contribute constructively to future states, akin to propagators in quantum field theory. Our proposed model enables efficient long-range dependency modeling while significantly reducing memory consumption compared to conventional Transformers.
In section V, we validate our approach through empirical evaluations, including the Passkey retrieval and summarization tasks. Our experimental results demonstrate that the proposed method achieves performance on par with the base model in contextual recall, while delivering superior memory efficiency compared to the base Llama-3.2 models [13].
@@ II Generalization of Transformer in Path Integral framework
Figure 1: Visualization of token state evolution in the Path Integral framework. The state $|x_{i},t_{n}\rangle$ at time $t_{n}$ evolves through multiple interaction pathways before reaching $|x_{i},t_{n+1}\rangle$(left side). Each evolution from $t_{n,\,j}$ to $t_{n,\,j}+\Delta t$ can be described as the transitions to all possible token states and the temporal evolutions of the projection states(right side).
The Path Integral approach provides a probabilistic framework for describing the evolution of a particle state by summing over all possible trajectories, with each path contributing according to a phase factor determined by the action. The foundation of this formulation lies in the Principle of Least Action, which states that the motion of a system follows a trajectory that minimizes (or extremizes) the action. In quantum mechanics, the Path Integral formulation extends the concept of the Least Action Principle by incorporating all possible paths leading to a future state, with their probability amplitudes interfering constructively or destructively.
Building upon this interpretation, we propose a novel generalization of the Transformer mechanism using the Path Integral formalism. In standard Transformers, the attention operation first determines the relationships between tokens within a sequence, followed by the Feed-Forward Network processing this contextualized information. This process bears a strong resemblance to the way quantum states evolve over time through the summation over all possible paths. By drawing on this analogy in the left side of Fig. 1, we reinterpret the transformation of a token vector through Transformer layers as the time evolution of a token state, where each token traverses multiple interaction pathways before contributing to the final output. Specifically, when we consider a case that a token state $|x_{i},\,t_{n}\rangle$ at position $x_{i}$ in a sequence and time $t_{n}$ temporally evolves to the state $|x_{i},\,t_{n+1}\rangle$ at position $x_{i}$ and time $t_{n+1}$, we can formulate this transition within a Path Integral framework as follows:
@@
where ${\cal L}$ is the Lagrangian of the Least Action Principle, ${\cal N}$ is overall normalization factor to the contribution of all possible paths and ${\cal T}$ is the operator of the time-ordered product. In the condition $\Delta t=\frac{t_{n+1}-t_{n}}{N}\rightarrow 0$, the equation (1) can be rewritten as
$|x_{i},\,t_{n+1}\rangle$ $={\cal N}\sum_{\text{all paths}}{\cal T}\Bigg{\{}\prod_{j=1}^{N}\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{n}+j\Delta t\right)\Delta t\right)\Bigg{\}}$
$\quad|x_{i},\,t_{n}\rangle\,.$ (2)
To consider the all possible paths in the way from $t_{n}$ to $t_{n+1}$, we input the identity matrix as
$I_{j}={\cal N}_{j}\sum_{k}|x_{k},\,t_{n,\,j}\rangle\langle x_{k},\,t_{n,\,j}|$ (3)
with the normalization factor ${\cal N}_{j}$ at each time $t_{n,\,j}\equiv t_{n}+\left(j-1\right)\Delta t$, which supposes transitions to the all possible token states at $t_{n,\,j}$. Here, we note that $x_{k}$ means $k$-th token position which is different from the ordinary position in space-time. Then, finally we get
$|x_{i},\,t_{n+1}\rangle$ $={\cal N}{\cal T}\bigg{\{}\prod_{j=1}^{N}\Bigg{[}\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{n}+j\Delta t\right)\Delta t\right)$
$\quad{\cal N}_{j}\sum_{k}|x_{k},\,t_{n,\,j}\rangle\langle x_{k},\,t_{n,\,j}|\Bigg{]}\bigg{\}}|x_{i},\,t_{n}\rangle\,.$ (4)
In particular, we take into account the time evolution of the state from $t_{n}+(j-1)\Delta t$ to $t_{n}+j\Delta t$, from the equation (4) as
$|x_{i},\,t_{n}+j\Delta t\rangle=\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{n}+j\Delta t\right)\Delta t\right)$
$\quad{\cal N}_{j}\sum_{k}|x_{k},\,t_{n,\,j}\rangle\langle x_{k},\,t_{n,\,j}|x_{i},\,t_{n,\,j}\rangle$ (5)
@@
$|x_{i},\,t_{n,\,j}\rangle$$\Leftrightarrow$ $Q^{T}=W_{Q}^{T}X_{i}^{T}$ (7)
$\langle x_{k},\,t_{n,\,j}|$$\Leftrightarrow$ $K=X_{k}W_{K}$ (8)
$|x_{k},\,t_{n,\,j}\rangle$$\Leftrightarrow$ $V^{T}=W_{V}^{T}X_{k}^{T}$ (9)
$\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{n}+j\Delta t\right)\Delta t\right)$$\Leftrightarrow$ $W_{FF}^{T}$ (10)
$|x_{i},\,t_{n}+j\Delta t\rangle$$\Leftrightarrow$ $X_{i}^{\prime\,T}\,,$ (11)
where $FF$ stands for Feed-Forward Network and $W_{Q,\,K,\,V}\in\mathbb{R}^{\,d_{E}\times d_{M}}$ are query, key and value matrices of the Transformer. That is the generalized form of decoder-only Transformer. Here we ignored multi-head attention for simplicity.
We introduced the concept of time into the Transformer structure by generalizing it within the Path Integral framework and establishing formal correspondences. This reinterpretation allows us to view the Transformer mechanism as the temporal evolution of an input token state. In this perspective, attention serves as a transition mechanism, enabling an input token to explore all possible token states within the same timeline. As the input token propagates through successive Transformer layers, its evolution follows a trajectory that encompasses all potential positional and temporal paths.
Training the Transformer can then be understood as an optimization process that identifies the Least Action path within this Path Integral formulation. Specifically, minimizing the cross-entropy loss between the predicted outputs and target tokens can be interpreted as selecting the optimal trajectory that satisfies the Principle of Least Action. Mathematically, this is expressed as
@@
Consequently, this framework provides an intuitive explanation for the superior performance of attention-based Transformers in next-token prediction. By dynamically refining token transitions through a learned path integral landscape, the Transformer effectively selects the most efficient computational paths that yield coherent and contextually relevant sequences.
In the next section, we will explore practical applications of this perspective in modeling time evolution within neural architectures.
@@ III Folded context Condensation
Figure 2: Underlying architecture of the model. A segment $|x^{s},\,t_{n}^{s}\rangle$ from the previous layer is projected into a query and also concatenated with time-evolved memory buffer $|x^{s-1},\,t_{n}^{s-1}+\Delta t\rangle$ to be key and value. The output of the current layer will be saved as a memory buffer and retrieved when the next segment arrives.
In our Transformer architectures, token states undergo temporal evolution across layers, leading to phase shifts that influence coherence in long-range dependencies. If unregulated, these shifts may induce destructive interference between token states of separate segments, causing information loss. To address this, we analyze the impact of token state phase shifts and introduce a periodic phase gap mechanism that ensures constructive interference.
As depicted in Fig. 2, the token state at position $i$ in segment $s$ evolves temporally through the Transformer layer as
$|x_{i}^{s},t_{n}^{s}+\Delta t\rangle$ $=\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{n}^{s}+\Delta t\right)\Delta t\right)$
$\quad{\cal N}\sum_{k}|x_{k}^{s},t_{n}^{s}\rangle\langle x_{k}^{s},t_{n}^{s}|x_{i}^{s},t_{n}^{s}\rangle\,.$ (14)
This equation suggests that each Transformer layer, after training, exhibits distinct phase evolution profiles aligning with eigenmodes of token state transitions. The layer parameters adapt to minimize phase misalignment, leading to the emergence of layer-specific eigen-frequencies.
This phenomenon is analogous to resonant systems such as stringed or wind instruments, where standing wave patterns form due to imposed boundary conditions. Consequently, the phase evolution of token states in a Transformer layer follows the wave equation:
@@
where $\Psi_{i}^{s}$ denotes the token state at position $i$ in segment $s$, and $\mathcal{H}$ represents the Hamiltonian of the layer’s transformation. This equation leads to standing wave solutions constrained by the boundary conditions between input and output tokens, indicating that token states evolve according to intrinsic phase shifts governed by learned parameters. These parameters are trained to align phase shifts with resonant frequencies, $f_{k}$’s, at each layer:
@@
where $\omega_{n}$ is the overall frequency, a function of resonant frequencies, and given by the input token state $\Psi_{i}^{s}$ applied by Hamiltonian. We note that $\Delta t_{n}$ is the time period of the standing wave at the $n$-th layer. After training, these values are fixed to ensure that token state transitions reinforce the context established by prior layers rather than interfere destructively.
To prevent destructive interference, we introduce a periodic phase gap between segments, enforcing coherence in token state transitions. Misaligned phase differences between adjacent segments can introduce interference effects, hampering information retention. We define the segment-to-segment transition as
$|x_{i}^{s},t_{n}^{s}+\Delta t\rangle$ $=\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{n}^{s}+\Delta t\right)\Delta t\right)$
$\quad{\cal N}\sum_{k}\left(\left|x_{k}^{s-1},t_{n}^{s-1}+\Delta t\rangle\langle x_{k}^{s-1},t_{n}^{s-1}+\Delta t\right|\right.$
@@ IV The implementation of the model
Following self-attention, the output undergoes residual connection processing and is passed through a Feed-Forward Network (FFN). The final layer-wise outputs align with corresponding segment-wise targets, yielding the total loss function:
@@
Figure 4: Each output of each layer for input segment is preserved as a memory buffer. Then, the memory buffer is concatenated with the current segment and takes part in the attention.

### v5

@@ I Introduction
Transformers architecture [1] has revolutionized natural language processing (NLP) by demonstrating unparalleled effectiveness in capturing long-range dependencies and enabling scalable, parallelized computation. This architecture often consists of an embedding layer, decoder layers with a self-attention mechanism, and a Feed-Forward Network (FFN), therefore, it predicts the next token in an auto-regressive manner. Unlike earlier Recurrent Neural Network (RNN) based models, which process sequences sequentially and suffer from vanishing gradient issues [2], Transformers leverage self-attention mechanisms to establish global dependencies across tokens, leading to significant improvements in contextual reasoning and coherent text generation. This architectural advantage has propelled Transformers to the forefront of modern NLP, underpinning the development of large-scale language models (LLMs) with remarkable generative capabilities.
However, Transformers have suffered from their computational inefficiency when handling long sequences. The self-attention mechanism inherently scales quadratically with sequence length in both memory and computation, making it impractical to process extensive contexts within a single pass. This limitation hinders the retention of dependencies across long documents, dialogues, or continuous interactions, prompting the development of alternative strategies such as Infini-Attention [7], which restructures the self-attention mechanism by introducing a hybrid memory system that allows interaction between local context tokens and external memory components, leveraging a linear attention [8] framework to reduce computational overhead. Another innovative solution is Titans [9], which incorporates trainable neural long-term memory, enabling the model to retain and retrieve past contextual information beyond traditional Transformer constraints.
Despite their empirical success, moreover, the theoretical foundations of Transformer architecture remain incomplete. While extensive research has explored their scaling properties [3], emergent behaviors [4], and empirical performance [5, 6], these studies provide limited insight into the fundamental mechanisms governing Transformer architectures. In particular, there is a lack of rigorous theoretical formulations explaining how information propagates through layers and how hierarchical structures are captured within Transformer models. This gap in understanding limits the interpretability and optimization of Transformer-based architectures and presents challenges in extending their capabilities systematically.
@@
Building on this theoretical foundation in Section II, we present an implementation in Section III and IV that extends Transformer architectures with a time-evolving memory buffer. This structured approach ensures that past token representations contribute constructively to future states, akin to propagators in quantum field theory. Our proposed model enables efficient long-range dependency modeling while significantly reducing memory consumption compared with conventional Transformers.
In section V, we validate our approach through empirical evaluations, including the Passkey retrieval and summarization tasks. Our experimental results demonstrate that the proposed method achieves performance on par with the base model in contextual recall while delivering superior memory efficiency compared with the base Llama-3.2 models [13].
@@ II Generalization of Transformer in Path Integral framework
Figure 1: Visualization of token state evolution in the Path Integral framework. The state $|x_{i},t_{0}\rangle$ at time $t_{0}$ evolves through multiple interaction pathways before reaching $|x_{i},t_{f}\rangle$(left side). Each evolution from $t_{j}$ to $t_{j}+\Delta t$ can be described as the transitions to all possible token states and the temporal evolutions of the projection states(right side).
The Path Integral approach provides a probabilistic framework for describing the evolution of a particle’s state by summing over all possible trajectories, with each path contributing according to a phase factor determined by the action. In classical mechanics —where the Planck constant $h$(=$2\pi\hbar$) is negligible with respect to the system’s characteristic energy scale— phase factors of the form $\exp(iS/\hbar)$ oscillate rapidly over time and cancel out for all but those paths near the classical trajectory. Consequently, only the trajectory that minimizes the action survives, so that the system’s motion follows the Principle of Least Action. Thus, the Path Integral formulation extends the Least Action Principle by incorporating every conceivable path to a future state, whose probability amplitudes interfere constructively or destructively. Here, the ratio of the system’s energy (or action) scale to the Planck constant determines whether its evolution appears essentially deterministic or fundamentally probabilistic.
Building upon this interpretation, we propose a novel generalization of the Transformer mechanism using the Path Integral formalism. In standard Transformers, the attention operation first determines the relationships between tokens within a sequence, followed by the FFN processing this contextualized information. This process bears a strong resemblance to the way quantum states evolve over time through the summation over all possible paths. By drawing on this analogy in the left side of Fig. 1, we reinterpret the transformation of a token vector through Transformer layers as the time evolution of a token state, where each token traverses multiple interaction pathways before contributing to the final output. Specifically, when we consider a case that a token state $|x_{i},\,t_{0}\rangle$ at position $x_{i}$ in a sequence and time $t_{0}$ temporally evolves to the state $|x_{i},\,t_{\rm f}\rangle$ at position $x_{i}$ and time $t_{\rm f}$, we can formulate this transition within a Path Integral framework as follows:
@@
where ${\cal L}$ is the Lagrangian governing the dynamics (analogous to the action functional in classical mechanics), ${\cal N}$ is overall normalization factor to the contribution of all possible paths and ${\cal T}$ is the operator of the time-ordered product. Here, it should be noted that $\hbar$ does not correspond to the physical value of the Planck constant, but rather functions as a tunable parameter reflecting the characteristic energy scale of the token system under consideration. In the condition $\Delta t=\frac{t_{\rm f}-t_{0}}{N}\rightarrow 0$, (1) can be rewritten as
$|x_{i},\,t_{\rm f}\rangle$ $={\cal N}\sum_{\text{all paths}}{\cal T}\Bigg{\{}\prod_{j=1}^{N}\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{0}+j\Delta t\right)\Delta t\right)\Bigg{\}}$
$\quad|x_{i},\,t_{0}\rangle\,.$ (2)
To consider the all possible paths in the way from $t_{0}$ to $t_{\rm f}$, we input the identity matrix as
$I_{j}={\cal N}_{j}\sum_{k}|x_{k},\,t_{j}\rangle\langle x_{k},\,t_{j}|$ (3)
with the normalization factor ${\cal N}_{j}$ at each time $t_{j}\equiv t_{0}+\left(j-1\right)\Delta t$, which supposes transitions to the all possible token states at $t_{j}$. Here, we note that $x_{k}$ means $k$-th token position which is different from the ordinary position in space-time. Then, finally we get
$|x_{i},\,t_{\rm f}\rangle$ $={\cal N}{\cal T}\bigg{\{}\prod_{j=1}^{N}\Bigg{[}\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{0}+j\Delta t\right)\Delta t\right)$
$\quad{\cal N}_{j}\sum_{k}|x_{k},\,t_{j}\rangle\langle x_{k},\,t_{j}|\Bigg{]}\bigg{\}}|x_{i},\,t_{0}\rangle\,.$ (4)
In particular, we take into account the time evolution of the state from $t_{0}+(j-1)\Delta t$ to $t_{0}+j\Delta t$, from (4) as
$|x_{i},\,t_{j+1}\rangle=\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{j}+\Delta t\right)\Delta t\right)$
$\quad{\cal N}_{j}\sum_{k}|x_{k},\,t_{j}\rangle\langle x_{k},\,t_{j}|x_{i},\,t_{j}\rangle$ (5)
@@
$|x_{i},\,t_{j}\rangle$$\Leftrightarrow$ $Q^{T}=W_{Q}^{T}X_{i}^{T}$ (7)
$\langle x_{k},\,t_{j}|$$\Leftrightarrow$ $K=X_{k}W_{K}$ (8)
$|x_{k},\,t_{j}\rangle$$\Leftrightarrow$ $V^{T}=W_{V}^{T}X_{k}^{T}$ (9)
$\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{j}+\Delta t\right)\Delta t\right)$$\Leftrightarrow$ $W_{FF}^{T}$ (10)
$|x_{i},\,t_{j+1}\rangle$$\Leftrightarrow$ $X_{i}^{\prime\,T}\,,$ (11)
where $FF$ stands for Feed-Forward Network and $W_{Q,\,K,\,V}\in\mathbb{R}^{\,d_{E}\times d_{M}}$ are query, key and value projection matrices. That is the generalized form of decoder-only Transformer. Here we ignored multi-head attention for simplicity.
Training the Transformer can then be understood as an optimization process that identifies the paths near the Least Action within this Path Integral formulation. Specifically, minimizing the cross-entropy loss between the predicted outputs and target tokens can be interpreted as selecting the optimal trajectories that satisfy the Principle of Least Action. Mathematically, this is expressed as
@@
Building on the above correspondence, we refine our interpretation of Transformers by viewing each Transformer layer as a temporal slice through which all possible computational paths evolve toward the final output state. Each layer, composed of an FFN and an attention mechanism, represents a discrete point in this evolution. A token state undergoes temporal evolution each time it passes through the FFN, resulting in a phase shift of the state. The attention mechanism, in contrast, evaluates the contextual relevance among all input token states. If the phase difference between two tokens leads to destructive interference, these token states are less likely to exchange contextual information. Therefore, the temporal evolution of a token —particularly the associated phase shift induced by the FFN— needs to be carefully treated in inter-token interactions. It is important to note that tokens located at different positions within a sequence do not necessarily correspond to distinct temporal moments. In this formulation, time is not defined by positional indices but emerges through the accumulated phase shifts governed by the FFN at each layer, leading to the final output. Accordingly, the primary factor to consider is the relative phase difference between token states situated at different positions.
In the following section, we explore the practical implications of this framework for modeling how contextual information can be preserved across segments that are input into Transformer layers in sequential orders.
@@ III Folded context Condensation
Figure 2: Underlying architecture of the model. A segment $|x^{s},\,t_{n}^{s}\rangle$ from the previous layer is projected into a query and also concatenated with time-evolved memory buffer $|x^{s-1},\,t_{n}^{s-1}+\Delta t_{n}\rangle$ to be key and value. The output of the current layer will be saved as a memory buffer and retrieved when the next segment arrives.
In our Transformer architectures, token states undergo temporal evolution governed by the FFN across layers, leading to phase shifts that influence coherence in long-range dependencies. If unregulated, these shifts may induce destructive interference between token states of separate segments, causing information loss. To address this, we analyze the impact of token state phase shifts and introduce a periodic phase gap mechanism that ensures constructive interference.
By rewriting (5), the token state at position $i$ in segment $s$ evolves temporally through the FFN at each Transformer layer as
$|x_{i}^{s},t_{n+1}^{s}\rangle$ $=\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{n}^{s}+\Delta t\right)\Delta t\right)$
$\quad{\cal N}\sum_{k}|x_{k}^{s},t_{n}^{s}\rangle\langle x_{k}^{s},t_{n}^{s}|x_{i}^{s},t_{n}^{s}\rangle\,,$ (14)
where we set the time of input segment $s$ into $n$-th layer as $t_{n}^{s}$. Here, we denote that the time of a token state represents the phase of the state, so the states sharing same time exhibit constructive or zero phase difference between them. This equation suggests that each Transformer layer, after training, exhibits distinct phase evolution profiles aligning with eigenmodes of token state transitions. Each layer parameters adapt to minimize phase misalignment, leading to the emergence of layer-specific eigen-frequencies.
This phenomenon is analogous to resonant systems such as stringed or wind instruments, where standing wave patterns form due to imposed boundary conditions. Consequently, the phase evolution of token states given in (14) follows the wave equation:
@@
where $\Psi_{i}^{s}$ denotes the token state at position $i$ in segment $s$, and $\mathcal{H}$ represents the Hamiltonian of the layer’s transformation governed by the Lagrangian. This equation leads to standing wave solutions constrained by the boundary conditions between input and output tokens, indicating that token states evolve according to intrinsic phase shifts governed by learned parameters. These parameters are trained to align phase shifts with resonant frequencies, $f_{k}$’s, at each layer:
@@
where $\omega_{n}$ is the overall frequency, a function of resonant frequencies, and given by the input token state $\Psi_{i}^{s}$ applied by the Hamiltonian. We note that $\Delta t_{n}$ is the time period of the standing wave at the $n$-th layer, and each layer has its own eigen time period. After training, each time period at each layer is fixed to ensure that token state transitions reinforce the context established by prior layers rather than interfere destructively.
To prevent destructive interference, we introduce a periodic phase gap between segments, enforcing coherence in token state transitions. Misaligned phase differences between adjacent segments can introduce interference effects, hampering information retention. We define the segment-to-segment transition getting from (14) as
$|x_{i}^{s},t_{n}^{s}+\Delta t_{n}\rangle$ $=\exp\left(\frac{i}{\hbar}{\cal L}\left(t_{n}^{s}+\Delta t\right)\Delta t\right)$
$\quad{\cal N}\sum_{k}\left(\left|x_{k}^{s-1},t_{n}^{s-1}+\Delta t_{n}\rangle\langle x_{k}^{s-1},t_{n}^{s-1}+\Delta t_{n}\right|\right.$
@@ IV The implementation of the model
Following self-attention, the output undergoes residual connection processing and is passed through an FFN. The final layer-wise outputs align with corresponding segment-wise targets, yielding the total loss function:
@@
Figure 4: Each output of each layer for the input segment is preserved as a memory buffer. Then, the memory buffer is concatenated with the current segment and takes part in the attention.